C S $ AD $100 $ 160 GDP $100 200 300 400 500 600 700 I $80 80 80 80 220 280 340 80 400 80 460 80 Complete the columns on S, aHi, I need some help with these. Thanks!
Answer #1
GDP C S = GDP - C I AD = C + I
100 100 100 - 100 = 0 80 100+80 = 180
200 160 200-160 = 40 80 160+80 = 240
300 220 300-220 = 80 80 220+80 = 300
400 280 400-280 = 120 80 360
500 340 500-340 = 160 80 420
600 400 600-400 = 200 80 480
700 460 700-460 = 240 80 540

1. C+S = GDP
So, S = GDP - C
And, C+I = AD

2. MPC = Change in C/Change in GDP = (160-100)/(200-100) = 60/100 = 0.6
MPS =  Change in S/Change in GDP = (40-0)/(200-100) = 40/100 = 0.4

3. APC = C/GDP = 400/600 = 0.67
APS = S/GDP = 200/600 = 0.33

4. Equilibrium GDP = $300 as equilibrium occurs when GDP = AD. They are both equal at the level of GDP = $300.

(Note: As HOMEWORKLIB's policy, 4 subparts are to be answered at a time.)